OpenStudy (loser66):

how to write 1+i and 1-i under the form of cos and sin? Please, help

OpenStudy (turingtest):

with a single angle?

OpenStudy (loser66):

my question is where I tag you. sorry for that. I am there for 2 hours without any reply

OpenStudy (loser66):

I have to break it down to get the answer

OpenStudy (loser66):

here is the link http://openstudy.com/study#/updates/5297740ae4b05ceaab567e29

OpenStudy (turingtest):

use\[e^{i\theta}\cos\theta+i\sin\theta\]assuming we can use two different values for theta, what value should theta take on to get a 1 out of this expression?